# Can Newtonian Gravity Produce Quantum Entanglement?

### v2

@@ Abstract
In this note, we consider specific scenarios of matter-gravity interaction based on the source theory to clarify the issue regarding the Gravitationally Induced Entanglement (GIE) protocol, which posits that the detection of quantum entanglement induced by a pure gravitational effect serves as a probe of quantum gravity. Thus, the GIE protocol implies that classical gravity cannot produce or mediate quantum entanglement, to which a recent paper Aziz and Howl (2025) shows the opposite. In all three scenarios of matter-gravity interactions, including mini-superspace, semiclassical gravity, and stochastic gravity approaches, our results agree with the GIE protocol and its implications. More precisely, we show that the mesoscopic quantum bodies with their parities of mass quadrupoles interacting via Newtonian gravity can get entangled only if the parity of the gravitational tidal field sourced by the quantum body is also quantized. We also point out a possible loophole that could seemingly violate the GIE protocol due to the truncation of higher-order terms in the perturbative expansion by orders of the Newton constant.
I Introduction and Conclusion
A recent paper in Nature Aziz and Howl (2025) claims that classical gravity can produce quantum entanglement through the exchange of virtual quantum matter, even though the mediator, the gravity field, is classical. This raises the debate over whether classical gravity can produce or mediate quantum entanglement. Especially in Marletto and Vedral (2025); Marletto et al. (2025), the authors object to the claim based on a dynamics-independent quantum information scheme, called the Gravitationally Induced Entanglement (GIE) protocol Marletto and Vedral (2017); Bose et al. (2017), which states that the detection of quantum entanglement involving only the pure gravity effect is a probe to quantum gravity. Therefore, the GIE protocol implies that the local classical mediator systems, such as classical gravity, cannot create quantum entanglement. The seemingly contradictory statements between Aziz and Howl (2025) and Marletto and Vedral (2017); Bose et al. (2017) deserve further clarification, which we address in this note. For some experimental proposals to realize the GIE protocol, see Marshman et al. (2020); Bose et al. (2022); Beckering Vinckers et al. (2023).
Though the GIE argument is sound, it is very general without attribution to specific exemplar gravity-matter models. This can create loopholes in some technical, detailed calculations, such as the one in Aziz and Howl (2025). To address such issues, in this note, we consider a specific quantum-matter and gravity model from the perspective of source theory. In this way, we will see that all the subtle issues regarding GIE become transparent, thus justifying its validity in the concrete contexts. Firstly, we define the quantum bodies as the mesoscopic superposed states of a $Z_{2}$ parity of some physical observables, such as the sign of an electric dipole or a mass quadrupole. The GIE protocol is to examine whether gravity can entangle a pair of such quantum bodies that are spatially separated. These mesoscopic quantum bodies for the GIE protocol play a similar role to the microscopic spin-$1/2$ particles in the Stern-Gerlach experiment. Secondly, we will generalize the source theory for the classical Newtonian gravity to three possible scenarios of dealing with the gravity interactions between quantum bodies: (I) the mini-superspace approach to quantum gravity by just quantizing the parity of the gravitational tidal field sourced by the quantum mass quadrupole; (II) the semiclassical gravity approach by sourcing the Newtonian potential by the expectation value of the mass quadrupole; (III) the stochastic gravity approach by appending the semiclassical gravity with a Langevin-type noise due to fluctuation of the parity of quantum mass quadrupole.
@@ I Introduction and Conclusion
By studying the unitarily evolved final state of the two quantum bodies dictated by the quadrupole-quadrupole interactions in the above three scenarios, we conclude that the quantum bodies can be entangled in scenario (I), but not in scenarios (II) and (III). Therefore, our result is consistent with the GIE protocol. Moreover, in scenarios (II) and (III), we point out a possible loophole based on the perturbative calculation, which will yield an artifact entanglement production due to the neglect of the subleading terms by truncating the expansion at a particular order of the Newton constant. This could be a cause of the disagreement between Aziz and Howl (2025) and the GIE ones Marletto and Vedral (2017); Bose et al. (2017).
II Mesoscopic Quantum Bodies and Their Preparation
@@
All considerations regarding the detection of quantumness in gravity are based on a setup originally proposed by Feynman during the 1957 Chapel Hill conference on gravity Dewitt and Rickles (2011) to prepare a massive object in a quantum superposition of two positions, which we will refer to as a quantum body. Physically, this mass superposition can be realized as distributed N00N states with spatial separation of two different modes located respectively at $\vec{x}_{L}$ and $\vec{x}_{R}$ Dowling (2008); Kim et al. (2025),
$|{\rm N00N}\rangle={1\over\sqrt{2}}\big(|N;\vec{x}_{L}\rangle|0;\vec{x}_{R}\rangle+e^{iN\theta}|0;\vec{x}_{L}\rangle|N;\vec{x}_{R}\rangle\big)\;,$ (1)
@@
where $N$ is either the photon or the atom numbers, depending on the experimental setup. This N00N state has recently been adopted to study its quantum decoherence by a black hole Danielson et al. (2022a, b, 2023); Biggs and Maldacena (2024).
@@
Moreover, if the N00N state also carries some electric charge, then by placing a charge of opposite sign at $\vec{x}_{M}=(\vec{x}_{L}-\vec{x}_{R})/2$, the combination forms a quantum body with an electric dipole with its direction in a superposition state. As there is no negative mass, it is impossible to produce the quantum mass dipole analogue. However, it is possible to form the mass quadrupole with its sign in the superposition state as shown in Fig. 1. Therefore, we will consider only the quantum mass quadrupoles and their mutual gravitational interactions in this note.
Formally, we can associate with such a quadrupole observable a mini-superspace operator
$\hat{Q}_{ij}=Q_{ij}\hat{Z}\;$, (2)
where $\hat{Z}$ is a $Z_{2}$ operator for measuring the sign of the quadrupole, satisfying
$\hat{Z}|\pm;Q_{ij}\rangle=\pm|\pm;Q_{ij}\rangle\;,\qquad{\rm with}\quad\langle\pm;Q_{ij}|\pm;Q_{ij}\rangle=1\;,\quad{\rm and}\quad\langle\mp;Q_{ij}|\pm;Q_{ij}\rangle=0\;,$ (3)
@@
where $\hat{Q}_{ij}$’s eigenstates $|\pm;Q_{ij}\rangle$’s correspond to the classical quadrupole moments with definite signs. We can then define a general superposed state of quadrupole signs as follows,
@@
The N00N state (1) is just a special case with $\theta=\pm{\pi\over 4}$, $\phi=0$. Note that
$\langle\theta,\phi;Q_{ij}|\hat{Q}_{ij}|\theta,\phi;Q_{ij}\rangle=Q_{ij}\langle\theta,\phi;Q_{ij}|\hat{Z}|\theta,\phi;Q_{ij}\rangle=Q_{ij}\cos 2\theta\;.$ (5)
This mini-superspace approach to quantum bodies can be thought of as the mesoscopic version of microscopic quantum spin-$1\over 2$}. Its simplicity helps elucidate the issue discussed in this note, much like the role of quantum spin-$1\over 2$} in Stern-Gerlach experiments in demonstrating quantum superposition.
Figure 1: A quantum body of mass quadrupole with its sign in the equal-weight superposed state.
@@ III Newtonian Gravity of Quantum Bodies
$\Phi(\vec{x})={G_{N}m\over r}+{G_{N}p_{i}x_{i}\over r^{3}}+{G_{N}Q_{ij}x_{i}x_{j}\over r^{5}}+\cdots\;,$ (6)
where $G_{N}$ is the Newton constant, $m$, $p_{i}$, and $Q_{ij}$ are respectively the mass moments of monopole, dipole, and quadrupole of a compact body. Since there is no negative mass, the mass dipole can be made to vanish by placing the origin at the center of mass. Moreover, the quadrupole moments are symmetric and traceless, i.e., $Q_{ij}=Q_{ji}$ and $\sum_{i=1}^{3}Q_{ii}=0$.
@@
Combining (6) and (7), we can obtain the interaction between two static quadrupole moments $Q_{ij}^{(1,2)}$ with the superscript labeling the quantum bodies,
$H_{QQ}^{c}=-{1\over 2}\Big[Q^{(1)}_{ij}\partial_{i}\partial_{j}\Phi^{(2)}\big|_{\vec{x}=\vec{x}_{2}}+Q^{(2)}_{ij}\partial_{i}\partial_{j}\Phi^{(1)}\big|_{\vec{x}=\vec{x}_{1}}\Big]=-{G_{N}\over r^{5}_{12}}Q_{ij}^{(1)}T_{ijkl}Q^{(2)}_{kl}\;$, (8)
@@
with $\hat{r}:={\vec{x}_{1}-\vec{x}_{2}\over|\vec{x}_{1}-\vec{x}_{2}|}$ and $r_{12}:=|\vec{x}_{1}-\vec{x}_{2}|$. This is the tidal interaction between two quadrupole moments, and can also be reproduced from the leading-order post-Newtonian approximation of Einstein gravity Binnington and Poisson (2009); Kol and Smolkin (2012); Biggs and Maldacena (2024); Lin and Mondal (2024), with a possibly different $T_{ijkl}$ due to the choice of gauge fixing. Moreover, the subleading PN terms will involve the motion of the compact bodies Goldberger (2007); Porto (2016), which we will not consider in this note.
@@ III.1 Mini-superspace approach
The first is to take the mini-superspace approach to quantum gravity, so that the induced Newtonian potential inherits the quantum nature of the sign associated with the quadrupole, i.e.,
@@
where $\hat{Z}^{(a)}$ is as defined in (2) and (3). This may not be a consistent treatment of quantum gravity. Still, it is analogous to the mini-superspace approach to quantum cosmology DeWitt (1967); Hartle and Hawking (1983); Vilenkin (1984), which involves quantizing only the scale factor, not the full metric.
Generalizing (8) to this mini-superspace case, the Hamiltonian for the interaction between two such quantum bodies can be found to be
@@ III.2 Semi-classical gravity approach
Another popular approach is the so-called semiclassical gravity DeWitt (1975); Parker (1968); Hawking (1975), in which matter, but not gravity, is quantized. Thus, the Newton potential is then sourced not by $\hat{Q}_{ij}$, but by its expectation value with respect to the N00N-like state. That is,
$\Phi^{{\rm semi},(a)}_{\rm quad}={G_{N}x_{i}x_{j}\over r^{5}}\langle\theta^{(a)},\phi^{(a)};Q^{(a)}_{ij}|\hat{Q}^{(a)}_{ij}|\theta^{(a)},\phi^{(a)};Q^{(a)}_{ij}\rangle={G_{N}Q^{(a)}_{ij}x_{i}x_{j}\over r^{5}}\cos 2\theta^{(a)}\;,\qquad a=1,2\;.$ (12)
Even though the Newton potential sourced by a quantum body is a classical quantity, this hybrid system exhibits bizarre behavior when a measurement collapses the quantum body, e.g., from $\theta={\pi\over 4}$ to $\theta=0$. This will then result in a sudden jump of the Newton potential due to the wavefunction collapse Penrose (1996). In general relativity, this implies that the spacetime metric can be singular.
Due to the classical nature of $\Phi^{\rm semi}_{\rm quad}$, the generailization of (8) in the semiclassical gravity becomes
$\hat{H}_{QQ}^{\rm semi}={1\over 2}H_{QQ}^{c}\big[(I^{(1)}\otimes\hat{Z}^{(2)})\cos 2\theta^{(1)}+(\hat{Z}^{(1)}\otimes I^{(2)})\cos 2\theta^{(2)}]\;,$ (13)
where $I^{(a)}$ is the identity operator for the $a$-th quantum body. Note that the quantum nature of $\hat{H}_{QQ}^{\rm semi}$ inherits from the acted quantum body, not from the source Newton gravity field. We also note that the N00N state (1) yields vanishing $\Phi^{{\rm semi},(a)}_{\rm quad}$ due to $\theta^{(q)}={\pi\over 4}$.
@@ III.3 Stochastic gravity approach
The third approach is stochastic gravity, an extension of semiclassical gravity that incorporates the stochastic noise $\xi_{\mu\nu}$ resulting from the quantum fluctuations of matter in the so-called Einstein-Langevin equation Hu and Sinha (1995); Hu and Verdaguer (2008); Martin and Verdaguer (1999),
@@
The gravity field plays the role of a Brownian particle in the Langevin dynamics, kicked by the Gaussian random noise $\xi_{\mu\nu}$ due to the leading-order influence functional Hu and Sinha (1995). To the Newtonian order, this implies that the Newton potential of (12) should be appended by a random noise arising from the sign fluctuation of the quantum body’s quadrupole. This noise can then contribute to the quadrupole-quadrupole interaction. This is equivalent to replacing $\theta^{(a)}$ in (12) and (13) by appending a white noise fluctuation
$\theta^{(a)}\rightarrow\tilde{\theta}^{(a)}=\theta^{(a)}+\delta\theta$ (15)
with $\delta\theta$ denote a Gaussian white noise ${\cal N}(0,1)$. Therefore, the quadrupole-quadrupole $\hat{H}_{QQ}^{\rm stoc}$ for the stochastic gravity still takes the form of (13) but replacing $\theta^{(a)}$ by $\tilde{\theta}^{(a)}$.
IV Entangling quantum bodies by Newtonian gravity ?

### v3

@@ Abstract
We investigate whether Newtonian gravity can generate quantum entanglement between mesoscopic quantum bodies modeled as superposed mass quadrupoles using three complementary approaches: mini-superspace, semiclassical gravity, and stochastic gravity. We systematically analyze gravitationally induced entanglement (GIE) mechanisms and the conditions under which they can arise. Our results support the GIE hypothesis by showing that the mini-superspace framework, which quantizes the parity of the gravitational tidal field, can entangle spatially separate quantum bodies. In contrast, the semiclassical and stochastic gravity models, in which the tidal gravitational field sourced by the quantum bodies remains classical, fail to entangle the final state. These findings clarify recent claims that classical gravity might induce entanglement, and reveal how perturbative treatments can lead to misleading conclusions.
I Introduction
The unification of quantum mechanics and general relativity remains one of the most profound challenges in fundamental physics. In the absence of direct empirical evidence for quantum gravity at the Planck scale, recent focus has shifted toward low-energy laboratory tests that can witness the non-classicality of the gravitational field. Building on a Gedanken-experiment originally proposed by Feynman at the 1957 Chapel Hill conference [DeWittRickles2011], recent table-top proposals [Marletto:2017kzi, Bose:2017nin]—often referred to as the Bose-Marletto-Vedral (BMV) mechanism—aim to verify the quantum nature of gravity [^1] For several other experimental proposals to realize the GIE protocol, see [Marshman:2019sne, Bose:2022uxe, BeckeringVinckers:2023kco]. These proposals utilize the dynamics-independent quantum information scheme, known as the gravitationally induced entanglement (GIE) protocol. Based on the local operations and classical communication (LOCC) protocol, the GIE protocol asserts that if gravity can mediate entanglement between two spatially separated massive quantum bodies, the gravitational field itself must possess non-classical features. Consequently, the prevailing consensus within the quantum information framework is that a local classical mediator cannot generate entanglement.
A recent study [Aziz:2025ypo] posits that classical gravity, when coupled with mesoscopic quantum matter, can indeed generate entanglement, despite the classical nature of the mediating field. This claim stands in stark contrast to the dynamical constraints argued in [Marletto:2025asw, Marletto:2025fpm], which maintain that the detection of gravity-mediated entanglement constitutes a definitive probe of quantum gravity  [^2] See [Nandi:2024jyf] for a theretical model to explore the quantum nature of gravity.. The apparent contradiction between the explicit calculations in [Aziz:2025ypo] and the GIE no-go theorems [Marletto:2017kzi, Bose:2017nin] indicates a critical need to bridge the gap between abstract information-theoretic principles and concrete field-theoretic models. While the GIE argument is robust in its generality, its lack of attribution to specific exemplar gravity-matter models leaves it vulnerable to interpretational loopholes arising from technical subtleties in perturbative calculations.
In this Letter, we address this issue by analyzing entanglement dynamics within a rigorous source-theory framework. Rather than relying on generic arguments, we construct a concrete model of gravity-matter interaction to examine the validity of the GIE protocol against the claims of classical-induced entanglement. We define our quantum bodies as mesoscopic superposed states characterized by the $Z_{2}$ parity of physical observables, specifically modeling them as mass quadrupoles. These mesoscopic quantum bodies for the GIE protocol play an analogous role to the microscopic spin-$1/2$ particles in the Stern-Gerlach experiment.
To systematically test the capacity of gravity to entangle these sources, we generalize the classical Newtonian source theory into three distinct interaction scenarios: (I) the mini-superspace approach to quantum gravity by just quantizing the parity of the gravitational tidal field sourced by the quantum mass quadrupole; (II) the semiclassical gravity approach by sourcing the Newtonian potential by the expectation value of the mass quadrupole; (III) the stochastic gravity approach by appending the semiclassical gravity with a Langevin-type noise due to fluctuation of the parity of quantum mass quadrupole.
@@ I Introduction
By studying the unitarily evolved final state of the two quantum bodies governed by quadrupole-quadrupole interactions in the three scenarios above, we conclude that the quantum bodies can be entangled in scenario (I) but not in scenarios (II) and (III). Therefore, our result is consistent with the GIE protocol. Moreover, in scenarios (II) and (III), we point out a possible loophole based on the perturbative calculation, which will yield an artifact entanglement production due to the neglect of the subleading terms by truncating the expansion at a particular order of the Newton constant. This could be a cause of the disagreement between [Aziz:2025ypo] and the GIE ones [Marletto:2017kzi, Bose:2017nin].
II Preparation of mesoscopic quantum bodies
@@
Following Feynman’s original proposal [DeWittRickles2011], a massive object may be prepared in a spatial superposition, which we call a "quantum body". Physically, this mass superposition can be realized as distributed N00N states with spatial separation of two different modes located respectively at $\vec{x}_{L}$ and $\vec{x}_{R}$ [Dowling:2008pbf, Kim:2025hbh]:
$|{\rm N00N}\rangle=\frac{1}{\sqrt{2}}\big(|N;\vec{x}_{L}\rangle|0;\vec{x}_{R}\rangle+e^{iN\theta}|0;\vec{x}_{L}\rangle|N;\vec{x}_{R}\rangle\big)\;,$ (1)
@@
where $N$ denotes the number of photons or atoms, depending on the experimental setup. Such states have recently been employed to study their quantum decoherence in the presence of a black hole [Danielson:2021egj, Danielson:2022tdw, Danielson:2022sga, Biggs:2024dgp].
@@
Moreover, if the N00N state also carries an electric charge, then by placing a charge of opposite sign at $\vec{x}_{M}=(\vec{x}_{L}-\vec{x}_{R})/2$, we can form a superposed electric dipole. In a gravitational analog—a mass dipole—is forbidden as there is no negative mass. However, it is possible to form the mass quadrupole with its sign in the superposition state as shown in Fig. 1(a) by superposing two proper arrangements of the relative positions of two mass dipoles. Consequently, we focus on constructing a quantum mass quadrupole in which the sign of the quadrupole moment is in superposition, and study the interactions of two such quantum bodies as shown in Fig. 1(b).
Figure 1: (a) Quantum body of mass quadrupole with its sign in the equal-weight superposed state. The $\pm\delta m$ are the local mass contrasts of either sign with respect to the underlying uniform mass distribution. (b) Two such quantum bodies with mass quadrupoles interact via the tidal gravitational fields. The main characteristics of this interaction are dictated by Newtonian gravity; however, its quantum nature will be scenario dependent.
@@
Formally, we can associate with such a quadrupole observable a mini-superspace operator $\hat{Q}_{ij}=Q_{ij}\hat{Z}$, where $Q_{ij}$ is a specific classical mass quadrupole, and $\hat{Z}$ is a $\mathbb{Z}_{2}$ operator for measuring the sign of the quadrupole, satisfying $\hat{Z}|\pm;Q_{ij}\rangle=\pm|\pm;Q_{ij}\rangle$ with $\langle\pm;Q_{ij}|\pm;Q_{ij}\rangle=1$ and $\langle\mp;Q_{ij}|\pm;Q_{ij}\rangle=0$. The $\hat{Q}_{ij}$’s eigenstates $|\pm;Q_{ij}\rangle$’s correspond to the classical quadrupole moments with definite signs. We can then define a general superposed state of quadrupole signs as follows,
@@
The N00N state (1) corresponds to the specific case where $\theta=\pm{\pi\over 4}$, $\phi=0$. The expectation value is
$\langle\theta,\phi;Q_{ij}|\hat{Q}_{ij}|\theta,\phi;Q_{ij}\rangle$ $=$ $Q_{ij}\langle\theta,\phi;Q_{ij}|\hat{Z}|\theta,\phi;Q_{ij}\rangle\;,$ (3)
$=$ $Q_{ij}\cos 2\theta\;.$
This mini-superspace approach to quantum bodies can be thought of as the mesoscopic version of microscopic quantum spin-$1\over 2$}. Its simplicity helps elucidate the issue discussed in this Letter, much as the role of the quantum spin-$1\over 2$} in Stern-Gerlach experiments demonstrates quantum superposition.
@@ III Newtonian Gravity of Quantum Bodies
$\Phi(\vec{x})=G_{N}\left({m\over r}+{p_{i}x_{i}\over r^{3}}+{Q_{ij}x_{i}x_{j}\over r^{5}}+\cdots\right)\;,$ (4)
where $G_{N}$ is the Newton constant, $m$, $p_{i}$, and $Q_{ij}$ are respectively the mass moments of monopole, dipole, and quadrupole of a compact body. This expansion is valid whenever the field distance $r$ is far larger than the size of the mass source, such as quantum bodies, which is the situation assumed in this Letter. Since there is no negative mass, the mass dipole can be made to vanish by placing the origin at the center of mass. Moreover, the quadrupole moments are symmetric and traceless, i.e., $Q_{ij}=Q_{ji}$ and $\sum_{i=1}^{3}Q_{ii}=0$.
@@
Combining (4) and (5), we can obtain the interaction between two static quadrupole moments $Q_{ij}^{(1,2)}$ depicted in Fig.1(b) with the superscript labeling the quantum bodies,
$H_{QQ}^{c}$ $=$ $-{1\over 2}\Big[Q^{(1)}_{ij}\partial_{i}\partial_{j}\Phi^{(2)}\big|_{\vec{x}=\vec{x}_{2}}+Q^{(2)}_{ij}\partial_{i}\partial_{j}\Phi^{(1)}\big|_{\vec{x}=\vec{x}_{1}}\Big]\;,$ (6)
$=$ $-{G_{N}\over r^{5}_{12}}Q_{ij}^{(1)}T_{ijkl}Q^{(2)}_{kl}\;$,
@@
with $\hat{r}:={\vec{x}_{1}-\vec{x}_{2}\over|\vec{x}_{1}-\vec{x}_{2}|}$ and $r_{12}:=|\vec{x}_{1}-\vec{x}_{2}|$. This is the tidal interaction between two static and well-separated quadrupole moments, and can also be reproduced from the leading-order post-Newtonian approximation of Einstein gravity [Binnington:2009bb, Kol:2011vg, Biggs:2024dgp, Lin:2024bmg], with a possibly different $T_{ijkl}$ due to the choice of gauge fixing. Moreover, the subleading PN terms will involve the motion of the compact bodies [Goldberger:2007hy, Porto:2016pyg], which we will not consider in this note.
@@ III.1 Mini-superspace approach
We first adopt a mini-superspace approximation to quantum gravity, wherein the induced Newtonian potential acquires a quantum character via the quadrupole sign operator:
@@
where $\hat{Z}^{(a)}$ represents the $\mathbb{Z}_{2}$ parity operator for the respective quantum bodies. While this may not be a consistent treatment of quantum gravity, it is conceptually analogous to the mini-superspace approach in quantum cosmology [DeWitt:1967yk, Hartle:1983ai, Vilenkin:1984wp], which involves quantizing only the scale factor, not the full metric.
Generalizing (6) to this framework, the Hamiltonian for the interaction between two such quantum bodies reads
@@ III.2 Semi-classical gravity approach
In the semi-classical gravity framework [dewitt1975quantum, parker1968particle, Hawking:1975vcx], matter is quantized while the gravitational field remains classical. Consequently, the Newtonian potential is sourced by the expectation value of the quadrupole operator with respect to the N00N-like state, rather than the operator itself:
$\Phi^{{\rm semi},(a)}_{\rm quad}$ $=$ $\frac{G_{N}x_{i}x_{j}}{r^{5}}\langle\theta^{(a)},\phi^{(a)}|\hat{Q}^{(a)}_{ij}|\theta^{(a)},\phi^{(a)}\rangle\;,$ (10)
$=$ $\frac{G_{N}Q^{(a)}_{ij}x_{i}x_{j}}{r^{5}}\cos 2\theta^{(a)}\;,\qquad a=1,2\;.$
Even though the Newton potential sourced by a quantum body is a classical quantity, this hybrid system exhibits bizarre behavior when a measurement collapses the quantum body, e.g., from $\theta={\pi\over 4}$ to $\theta=0$. This will then result in a sudden jump of the Newton potential due to the wavefunction collapse [Penrose:1996cv]. In general relativity, this implies that the spacetime metric can be singular.
Given the classical nature of the source in (10), the interaction Hamiltonian takes the form:
$\hat{H}_{QQ}^{\rm semi}=\frac{1}{2}H_{QQ}^{c}\left[(I^{(1)}\otimes\hat{Z}^{(2)})\cos 2\theta^{(1)}+(\hat{Z}^{(1)}\otimes I^{(2)})\cos 2\theta^{(2)}\right]\;,$ (11)
where $I^{(a)}$ denotes the identity operator for the $a$-th body. Crucially, the quantum nature of $\hat{H}_{QQ}^{\rm semi}$ arises solely from the probe body, not the source field. For the specific N00N state in (1), $\Phi^{{\rm semi},(a)}_{\rm quad}$ vanishes as $\theta^{(a)}=\pi/4$.
@@ III.3 Stochastic gravity approach
Stochastic gravity extends the semi-classical approximation by incorporating fluctuations in the matter stress-energy tensor via the Einstein-Langevin equation [Hu:1994dka, Hu:2008rga, Martin:1998nc]:
@@
Here, the gravitational field is driven by a Gaussian stochastic source $\xi_{\mu\nu}$, arising from the noise kernel (fluctuations of the quantum matter) [Hu:1994dka]. This is the generalization of Langevin dynamics for Brownian motion, with the gravity field as the Brownian particle and the quantum matter as the thermally-agitated water molecules. At the Newtonian order, the potential in (10) acquires a random component corresponding to fluctuations in the quadrupole orientation. This is modeled by modifying the orientation parameter $\theta^{(a)}$:
$\theta^{(a)}\rightarrow\tilde{\theta}^{(a)}=\theta^{(a)}+\delta\theta\;,$ (13)
where $\delta\theta$ represents Gaussian white noise $\mathcal{N}(0,1)$. Consequently, the stochastic interaction Hamiltonian $\hat{H}_{QQ}^{\rm stoc}$ retains the form of Eq. (11), with $\theta^{(a)}$ replaced by the stochastic variable $\tilde{\theta}^{(a)}$.
IV Does Newtonian gravity entangle quantum bodies?